What are the solutions to the system.
1. y = x² + 3x - 5
y = x + 3

Answer by htmentor(1343)

Solve by substitution:
x + 3 = x² + 3x - 5
Solving for x, we get x² + 2x - 8 = 0
This factors as (x-2)(x+4) = 0, which has solutions x=2, x=-4
Thus, y = 2+3 = 5, and y = -4 + 3 = -1
Ans: (2,5) and (-4,-1)